OpenStudy (anonymous):

Is this statement true or false? In a 45°-45°-90° triangle, the hypotenuse is 2 times as long as the leg.

OpenStudy (anonymous):

I'll fan! I need help!!

OpenStudy (anonymous):

If I remember correctly, Hypotenuse for a 45 45 90= \[\sqrt{2}Long leg\]

OpenStudy (anonymous):

so false

OpenStudy (ranga):



OpenStudy (skullpatrol):

$$\Huge a^2+b^2=c^2$$

Directrix (directrix):

See attachment for 45-45-90 theorem "refresher."